Circular No. 7649

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2001cq, 2001cr, 2001cs, 2001ct, 2001cu, 2001cv, 2001cw

M. Doi, Institute of Astronomy, University of Tokyo, reports the discovery of seven apparent supernovae by the Subaru high-redshift supernova search group (M. Doi, H. Furusawa, F. Nakata, M. Ouchi, N. Yasuda, S. Miyazaki, N. Kashikawa, Y. Komiyama, Y. Ohyama, M. Yagi, K. Aoki, I. Hook, S. Perlmutter, and G. Aldering), which is part of the Supernova Cosmology Project (cf. IAUC 6804). The 1-hr discovery exposures taken with the Subaru f/1.86 telescope (+ prime-focus Suprime-Cam) at Mauna Kea were subtracted from 1-hr exposures taken on Apr. 25; optical spectroscopy (range 470–920 nm) with Subaru’s FOCAS spectrograph was obtained on May 26 and 27 (exposures 10–105 min). The magnitudes are in the AB system (cf. Oke and Gunn 1983, Ap.J. 266, 713; Fukugita et al. 1996, A.J. 111, 1748; Fukugita et al. 1995, PASP 107, 954). The magnitudes for SNe 2001cq and 2001ct include their host galaxies. The spectra of SNe 2001cq, 2001cr, 2001ct, and 2001cu show broad features, though supernova types could not be determined (redshifts were estimated from the spectra). Each supernova is close to the center of its faint host galaxy (the host for SN 2001cr was not seen). Finding charts are available at

http://indus.astron.s.u-tokyo.ac.jp/works/suprime/SNe.

SN  2001 UT  α_{2000}  δ_{2000}  Mag.  z Type
2001cq May 19.6 16^{h}05^{m}18.^{s}70 +43^{o}11′42′′.1 22.8 R 0.20? Ia?
2001cr May 20.4 13 24 22.71 +27 34 33.9 23.8 I 1.02? Ia?
2001cs May 20.4 13 25 17.32 +27 13 24.2 23.5 I 0.41 Ia
2001ct May 20.4 13 24 44.63 +27 15 27.2 23.4 I 0.45? II?
2001cu May 20.4 13 25 32.73 +27 41 34.4 23.7 I 0.50? Ia?
2001cv May 20.4 13 25 34.84 +27 37 02.6 24.1 I 1.04 Ia
2001cw May 20.5 15 23 06.31 +29 39 31.9 24.3 I 0.93 Ia

SUPERNOVA 2001cp IN UGC 10738 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2001cp, obtained by M. Calkins on June 21.44 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova before maximum light. Adopting the NASA/IPAC Extragalactic Database recession velocity of 6716 km/s for the host galaxy, the supernova expansion velocity is ∼ 11 400 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 7 ± 2 days before maximum light.

2001 June 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green